A jet plane flying with the wind went 1030 mi in 2 h. Against the wind, the plane could only fly 820 mi in the same amount of time. Find the rate of the plane in calm air and the rate of the wind. (Round your answers to one decimal place.)
rate of plane ________ mph
rate of wind _________ mph

Found 2 solutions by mananth, richwmiller:
Answer by mananth(16949)

A jet plane flying with the wind went 1030 mi in 2 h. Against the wind, the plane could only fly 820 mi in the same amount of time. Find the rate of the plane in calm air and the rate of the wind. (Round your answers to one decimal place.)

let the speed of the plane in still air be p
the speed of the wind be w
2*(p+w)= 1030
2*(p-w)=820
p+w=510
p-w=410
2p=920
p= 460 mph
w= 510-460=50mph

Answer by richwmiller(17219)

2(p+w)=1030
2(p-w)=820
p+w=515 this is where the other tutor went wrong
p-w=410
2p=925
p=462.5 mph
462.5+w=515
w=52.5 mph